OpenStudy (anonymous):

a body released from the top of a tower of height h takes t second to reach the ground at time 1/2t the height of the body from the ground?

OpenStudy (anonymous):

Common,. this is easy , you got to give it a try yourself .

OpenStudy (anonymous):

*comm'on

OpenStudy (ajprincess):

use s=ut+1/2at^2 instead of t use t/2

OpenStudy (anonymous):

@ajprincess , let him think. This is one of the easy questions :)

OpenStudy (ajprincess):

ok @shivam_Bhalla

OpenStudy (anonymous):

thanxx for ur hints..

OpenStudy (anonymous):

i got the answer in terms of u the answer is 3h/4

OpenStudy (anonymous):

What is u in this case @shameer1 ??

OpenStudy (anonymous):

u=0

OpenStudy (anonymous):

i forgot it

OpenStudy (ajprincess):

nw wat do u get?

OpenStudy (anonymous):

5/4

OpenStudy (ajprincess):

check again.

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

did u check?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

r u sure of instead of t use t/2

OpenStudy (ajprincess):

ya aftr finding the answer in terms of h, subtract it frm h to get the final answer.

OpenStudy (anonymous):

h=1/2*10*t^2/4

OpenStudy (ajprincess):

here it's nt h bt sumthing lyk h1 or h'.

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

use s=ut+1/2at^2 again this time putting h for s nd t for time.